OpenStudy (anonymous):

if the volume of CYLINDER WITH A height of 3 feet is 75 pie cubic feet, find the surface area of the cylinder in square feet

OpenStudy (polaris_s0i):

first: the area of the cylinder is based on 2 things: height and radius. its the area of the circle, multiplied by the height. \[ \pi r^2h\]

OpenStudy (polaris_s0i):

you can reverse this to find the radius: \[ A = \pi r^2h \] \[ \frac{A}{\pi h} = r^2\] \[r=\sqrt{\frac{A}{\pi h}}\]

OpenStudy (polaris_s0i):

sorry that should be V in the equation above not A

OpenStudy (polaris_s0i):

that's for volume.

OpenStudy (polaris_s0i):

so \[ r = \sqrt{\frac{V}{\pi h}}\]

OpenStudy (polaris_s0i):

you can then use this information to find the surface area of the cylinder. it has the area of 2 circles at the top and bottom, and then the perimeter of a circle multiplied by the height. \[ SA = 2*\pi r^2 + 2\pi rh\] or \[SA = 2\pi r(r + h)\]

OpenStudy (polaris_s0i):

after doing some algebra, you could go directly to the surface area: \[SA = 2\pi\left(\sqrt{\frac{V}{\pi h}}\right)^2 + 2\pi h\sqrt{\frac{V}{\pi h}}\] \[SA = \frac{2\cancel{\pi} V}{\cancel{\pi} h} + \sqrt{\frac{4\pi^{\cancel{2}}h^{\cancel{2}}V}{\cancel{\pi} \cancel{h}}}\] and you're left with: \[SA = \frac{2V}{h} + \sqrt{4\pi hV}\]

OpenStudy (polaris_s0i):

\[SA = \frac{2*75\pi}{3} + \sqrt{4\pi*3*75\pi}\] \[SA = 50\pi + \sqrt{900\pi^2}\] \[SA = 50\pi + 30\pi\] \[SA = 80\pi\]